OpenStudy (turbostudier):

can someone give me a hand? (Monomials) 4abc^7 + 10a^3b^2c^5 if a=2 b=3 c=1, what do i get?

OpenStudy (turbostudier):

if 1 response will medal 2 responses, will medal 1, fan the other

OpenStudy (turbostudier):

plz help

jhonyy9 (jhonyy9):

so than given that a=2 ,b = 3 and c=1 just substitute these values in place of every term and calculi what will get

jhonyy9 (jhonyy9):

in this way 4*2*3*1^7 + ...

OpenStudy (turbostudier):

so 4abc^7.. i multiply them?

OpenStudy (turbostudier):

oh

jhonyy9 (jhonyy9):

yes there between these terms is multiplication

OpenStudy (turbostudier):

ohh ok just checking

OpenStudy (turbostudier):

wait so i get 24^7+20^33^31^5

jhonyy9 (jhonyy9):

24^7 how ? there is just 1^7

OpenStudy (turbostudier):

but u said i hve to multiply them

jhonyy9 (jhonyy9):

4*2*3*1^7 = 4*2*3*1 = 4*6= 24 and the other one

OpenStudy (phi):

c^7 or 1^7 means 1*1*1*1*1*1*1

OpenStudy (turbostudier):

so 4 times 2 times 3 times 1

jhonyy9 (jhonyy9):

yes

OpenStudy (turbostudier):

then to the ^7

jhonyy9 (jhonyy9):

just 1 is on exponent of 7

OpenStudy (phi):

do you remember order of operations you do the exponents first 1^7 becomes just 1 (because 1 *1 *1 *..1 is just 1)

jhonyy9 (jhonyy9):

or not ?

OpenStudy (turbostudier):

but i plugged in 2,3, and 1 correct? but i still had a 4 at the front..

jhonyy9 (jhonyy9):

yes

OpenStudy (turbostudier):

4abc^7

jhonyy9 (jhonyy9):

and 4*2*3*1^7 = 4*2*3*1 = ?

OpenStudy (turbostudier):

so are you saying i multiply 24 7 times??

jhonyy9 (jhonyy9):

no there is just 1 on exponent of 7

jhonyy9 (jhonyy9):

and you know that 1 on exponent of 7 = 1

OpenStudy (turbostudier):

oh so it becomes just 24?

jhonyy9 (jhonyy9):

yes sure and the other one than you substitute theses values ?

OpenStudy (turbostudier):

20^33^31^5 but it should be just 20^33^3

OpenStudy (turbostudier):

right?

jhonyy9 (jhonyy9):

10*2^3 *3^2 *1^5 = ?

jhonyy9 (jhonyy9):

2^3 = ?

OpenStudy (turbostudier):

20^3 * 3^2 * 1^5 i dont multiply tho i leave it like that

jhonyy9 (jhonyy9):

no

OpenStudy (turbostudier):

except for 1^5 disappears

jhonyy9 (jhonyy9):

there is just 2^3

OpenStudy (turbostudier):

how?

jhonyy9 (jhonyy9):

what how ? there is just 2 on exponent of 3 and 3 on exponent of 2 and 1 on exponent of 5

jhonyy9 (jhonyy9):

10 there have not exponent

OpenStudy (turbostudier):

yah

OpenStudy (phi):

turbo, you are not following the "order of operations" you start with 10a^3b^2c^5 now remember that is short for 10 * a^3 * b^2 * c^5 (we usually don't show the multiply signs, but they are there) then you replace the letters with the numbers they gave you a=2 b=3 c=1 10 * 2^3 * 3^2 * 1^7 now order of operations (people use PEDMAS or BODMAS to remember) but the idea is first do the exponents for example what is 2^3 ?

OpenStudy (turbostudier):

nmi think i got it

OpenStudy (turbostudier):

i am working it out pen and paper

OpenStudy (turbostudier):

so jhonyy9 explained that i get 24 for the first part

OpenStudy (phi):

yes

OpenStudy (turbostudier):

that means i get 20^3 and 3^3 for the second part

OpenStudy (phi):

to get that from 10 * 2^3 * 3^2 * 1^7 you must have done 10*2 first and that is not allowed we do exponents before multiply in other words, you have to first do 2^3 and 3^2 and 1^7

OpenStudy (turbostudier):

ohhh i see what u are saying.. i am very sorry

OpenStudy (turbostudier):

ok let me solve

OpenStudy (turbostudier):

720???????

OpenStudy (phi):

yes, that sounds right now add the first term (the 24) to get 720+24

OpenStudy (turbostudier):

222=8 33=9 1111111=1 *10= 720

OpenStudy (turbostudier):

ok

OpenStudy (turbostudier):

744

OpenStudy (phi):

but we would write 2^3 as 2*2*2 not 222 because 2*2*2 is 8 and that is much different from two hundred twenty two (222) yes, 744 is the answer

OpenStudy (turbostudier):

but im still a little confused for as of whythe first part is 24 tho 4abc^7 4*2*3*1 ^7

OpenStudy (phi):

ok, in that mess, which part do you do first ?

OpenStudy (turbostudier):

oh yah the OHHHHHXD

OpenStudy (turbostudier):

THE EXPONENT

OpenStudy (turbostudier):

ok that makes sense

OpenStudy (turbostudier):

i get,,,, 24+720

OpenStudy (turbostudier):

then i get.... =744

OpenStudy (phi):

yes

OpenStudy (turbostudier):

ok Thank you so much!!!